OpenStudy (anonymous):

CAN SOMEONE HELP ME ON THIS! Turner’s treadmill starts with a velocity of −2.7 m/s and speeds up at regular intervals during a half-hour workout.After 33 min, the treadmill has a velocity of −7.1 m/s. What is the average acceleration of the treadmill during this period? Answer in units of m/s

OpenStudy (anonymous):

\[\frac{V _{f}-V_{i}}{t}\]There's your formula. Convert your minutes to seconds, plug in your values, and you're done.

OpenStudy (amistre64):

is should be the slope between the points: (0,-2.7) and (33,-7.1) if i read it right

OpenStudy (amistre64):

which comes to what brandon posted :)

OpenStudy (anonymous):

whats the answer so that i know if i have the answer right.

OpenStudy (amistre64):

what answer did you get?

OpenStudy (amistre64):

\[30\ min*\frac{60\ sec}{1\ min}\] \[30\ \cancel{min}*\frac{60\ sec}{1\ \cancel{min}}\] \[30*60\ sec=1800\ sec\]

OpenStudy (amistre64):

correction; 33 60 --- 1980 33*60sec = 1980 sec for comparison

OpenStudy (anonymous):

I got 0.0022222?

OpenStudy (amistre64):

i get: (1980,-7.2) -( 0 ,-2.7) ------------- 1980, -4.5 slope = -4.5/1980 = -21/(5)1980 = -7/3300 or, -.002121....

OpenStudy (amistre64):

or something thereabouts ....

OpenStudy (amistre64):

are those spose to be negative signs or just a way to indicate with the number is?